### v1

# RomanLens: Latent Romanization and its role in Multilinguality in LLMs

Abstract
Large Language Models (LLMs) exhibit remarkable multilingual generalization despite being predominantly trained on English-centric corpora. A fundamental question arises: how do LLMs achieve such robust multilingual capabilities? For non-Latin script languages, we investigate the role of romanization—the representation of non-Latin scripts using Latin characters—as a bridge in multilingual processing. Using mechanistic interpretability techniques, we analyze next-token generation and find that intermediate layers frequently represent target words in romanized form before transitioning to native script, a phenomenon we term Latent Romanization. Further, through activation patching experiments, we demonstrate that LLMs encode semantic concepts similarly across native and romanized scripts, suggesting a shared underlying representation. Additionally in translation towards non Latin languages, our findings reveal that when the target language is in romanized form, its representations emerge earlier in the model’s layers compared to native script. These insights contribute to a deeper understanding of multilingual representation in LLMs and highlight the implicit role of romanization in facilitating language transfer. Our work provides new directions for potentially improving multilingual language modeling and interpretability.
RomanLens: Latent Romanization and its role in Multilinguality in LLMs
Alan Saji1222Work done during employment at AI4Bharat.333
Correspondence: Alan Saji (alansaji2001@gmail.com)
, Ratish Puduppully
(rapu@itu.dk)
Jaavid Aktar Husain2222Work done during employment at AI4Bharat.,
Thanmay Jayakumar1,3,
Raj Dabre1,3,4,5
Anoop Kunchukuttan1,3,6,
Mitesh M. Khapra1,3,
Ratish Puduppully7333
Correspondence: Alan Saji (alansaji2001@gmail.com)
, Ratish Puduppully
(rapu@itu.dk)
1Nilekani Centre at AI4Bharat,
2Singapore University of Technology and Design,
3Indian Institute of Technology Madras, India,
4National Institute of Information and Communications Technology, Kyoto, Japan,
5Indian Institute of Technology Bombay, India,
6Microsoft, India,
7IT University of Copenhagen
1 Introduction
The majority of modern Large Language Models (LLMs) Touvron et al. (2023); Dubey et al. (2024); Team et al. (2024) are trained predominantly on English-dominated corpora. Yet, they exhibit strong multilingual generalization across diverse languages Huang et al. (2023); Zhao et al. (2024a); Zhang et al. (2023). This raises a fundamental question: How do LLMs develop such robust multilingual capabilities despite their English-centric training?
To address this, we adopt a mechanistic interpretability perspective. Wendler et al. (2024) suggest that LLMs encode multilingual information within a shared, language-agnostic latent space, albeit with an inherent bias toward English due to training data composition and architectural choices. Further, Kojima et al. (2024) describe distinct phases in multilingual information processing: initial layers map language-specific lexical and syntactic representations to a language-independent semantic space, middle layers maintain this semantic abstraction, and final layers transform these representations into language-specific lexical and syntactic forms.
In this study, we focus on languages written in non-Latin scripts. We hypothesize that LLMs leverage romanized forms of non-Latin script languages as an intermediate bridge between their language-agnostic concept space and language-specific output representations. Romanization—the representation of non-Latin scripts using Latin characters—may facilitate this process by aligning non-English languages more closely with English. Supporting this, Husain et al. (2024) demonstrated that explicitly romanizing inputs improves model performance on multilingual tasks, suggesting an inherent alignment between romanized text and English representations. We investigate whether LLMs indeed use romanization as a bridge between language-agnostic concepts and language-specific outputs given its potential implications for understanding multilingual processing in LLMs.
[…]
To further probe this phenomenon, we employ activation patching Ghandeharioun et al. (2024); Variengien and Winsor (2023); Chen et al. (2024); Dumas et al. (2024), a technique that replaces activations from one forward pass with another to analyze the resulting outputs (see Section 3.2). Dumas et al. (2024) found that LLMs process language and conceptual information as distinct entities. Building on this, we perform layerwise activation patching between romanized and native-script inputs to examine whether LLMs encode conceptual information similarly across scripts.
, ’pa’ is its romanized representation). Note that push is a prefix of pushpam, the romanized form of the translation of “flower”; also, poo is another romanized translation of “flower.”
Based on our experiments, we list down a summary of our contributions below:
1.
Latent Romanization Across Layers: During multilingual next-token generation, intermediate layers occasionally represent tokens in romanized form before resolving to native script (Figure 1). We term this phenomenon Latent Romanization.
2.
Consistent Semantic Encoding Across Scripts: Activation patching experiments reveal that LLMs encode semantic concepts similarly, regardless of whether the input is in native or romanized script.
3.
Earlier Emergence of Target Representations: When translating into romanized versus native script, romanized target representations emerge earlier in the model’s layers—typically one or two layers prior to native script representations (Figure 2). This suggests that romanization facilitates faster progression toward language-specific embeddings.
2 Related Work
Recent studies investigate latent representations in multilingual tasks, examining whether English emerges as a latent language in English-centric LLMs Wendler et al. (2024) and how the composition of training corpus mixtures influences latent representations Zhong et al. (2024). Recent investigations analyze how LLMs handle multilingual capabilitiesZhao et al. (2024b). Interpretability tools relevant to this work include logit lens Nostalgebraist (2020), tuned lens Belrose et al. (2023) and direct logit attribution Elhage et al. (2021) which are key tools for decoding intermediate token representations in transformer models. The logit lens applies the language modeling head to earlier layers without additional training, while the tuned lens in addition to this trains an affine mapping to align intermediate states with final token predictions. Direct logit attribution attributes logits to individual attention heads. This work focuses on the logit lens (Section 4.1) to investigate whether English-centric decoder only LLMs like Gemma-2, when prompted in non-latin language, processes via romanized latent states before producing non-English text. Tuned lens is avoided as its training process might obscure the intermediate romanized states by aligning them to final non-English outputs, potentially masking the phenomenon under investigation.
Activation patching (Meng et al., 2022) is a key interpretability technique employed in our study. This technique has been used to draw causal interpretations of LLMs representations Variengien and Winsor (2023); Geiger et al. (2022); Kramár et al. (2024); Ghandeharioun et al. (2024); Chen et al. (2024). Building on these approaches, we adopt an activation patching-based experimental framework as defined in Dumas et al. (2024) to investigate and compare how concepts are encoded in romanized versus native scripts.
Previous studies have demonstrated that romanization can serve as an effective approach to interact with LLMs Husain et al. (2024). Liu et al. (2024) and Xhelili et al. (2024) employ an approach based on contrastive learning for post-training alignment, contrasting sentences with their transliterations in Latin script to overcome the script barrier and enhance cross-lingual transfer.
However, our work distinguishes itself from prior research by exploring the presence of romanized representations in the latent layers of an LLM during multilingual tasks, an aspect that, to the best of our knowledge, has not yet been investigated.
3 Background
We give a quick background of the transformer’s forward pass as well as the basics of mechanistic interpretability approaches such as logit lens and activation patching which we leverage in this paper.
Transformer’s forward pass: Here, we study autoregressive decoder-only transformer models. Vaswani (2017). Such models maintain the dimensionality of the input data throughout the computation process. For an input sequence , where is the sequence length, each token is initially mapped to an embedding vector, which is then transformed through the model’s layers, producing a sequence of hidden states (also referred to as latents).
The initial latents are obtained from a learned embedding matrix, which maps each vocabulary token to a fixed vector. These latents are updated iteratively across layers through the addition of residuals. The update rule for the latent at position in layer is expressed as:
The function , , is known as a transformer block, where is the final layer index. For prediction, the final embedding vector is multiplied by an unembedding matrix , producing a vector containing the logit scores for each vocabulary token . This is computed as:
These logit scores are converted to probabilities via the softmax function as follows:
3.1 Interpretability tool: Logit lens
Generally, in a decoder only LLM, the unembedding matrix is multiplied with the final hidden state and a softmax is taken on the product to produce the token distributions at that token generation step. Since all hidden states of an LLM are in the same shape, it is possible to apply the unembedding matrix and softmax on all layers, thereby generating token distributions at all layers. This method of prematurely decoding hidden states is referred to as logit lens Nostalgebraist (2020). Logit lens reveals how the latent representations evolve across layers to produce the final output, providing insights into the progression of computations within the model.
3.2 Interpretability tool: Activation patching
Activation patching involves modifying or patching the activations at specific layers during a forward pass and observing the effects on the model’s output. In this work, we adopt the activation patching setup introduced in Dumas et al. (2024).
In the context of activation patching, let denote the language of a word, denote the concept of a word and denote that word. Eg: If = cow and = ‘en’ then = ‘cow’. Similarly = ‘vache’. We use 5-shots translation prompts to create paired source and target prompt , with different concept, input language, and output language. Unless otherwise specified, and refer to the output languages of and , respectively.
[…]
4 Methodology
We design our analysis setup with the intention of addressing the following research questions:
RQ1: Does LLMs exhibit latent romanization during multilingual text completion tasks? (Section 4.1)
RQ2: What are the differences in hidden layer representations when processing the same language in romanized and native scripts? (Section 4.2)
RQ3: How does the representation of semantic concepts in LLMs compare between native and romanized scripts of non-Latin languages? (Section 4.3)
4.1 Logit lens
Prompt design. In order to have the next token be easily inferred from the given prompt , we establish the following prompts.We use Hindi as the example language to describe the prompts. English translation and transliteration of the below prompts are present in Appendix A.
[…]
Repetition task. We ask the model to repeat the word mentioned in the same language.
[…]
Cloze task. We ask the model to predict the masked word in a sentence. We give the model two examples and ask it to predict the masked word in the next masked sentence.
[…]
Latent Romanization Analysis
Translation, repetition, and cloze tasks are explored by providing the respective prompts as inputs to an LLM to generate the corresponding output word. We romanize the output word, tokenize it, retaining only the tokens present in the model’s vocabulary, and analyze the occurrence of these tokens in the latent layers across timesteps of the output word generation. The analysis is done using logit lens by examining whether the probability of a romanized token in the next token distribution at a given layer exceeds 0.1 - a threshold optimized for detection accuracy i.e minimizing false positives and maximizing true positives. We refer to this hereafter as the latent romanization condition. Our analysis focuses on the final 10 layers of an LLM, where coherent romanized representations emerge according to logit lens visualizations (c.f. Figures 1 and 9 - 13).
We track romanized tokens using a timestep-specific tokenization scheme optimized for detection accuracy. In the first output generation timestep, we check for tokens that include the full romanized word and its prefixes. During intermediate timesteps,we check for all possible substrings of the romanized word in the latent layers. In the final output generation timestep, we probe the presence of only the full romanized word and its suffixes as potential tokens (c.f. Appendix B).
Latent romanization is analyzed under three distinct scenarios for the last 10 layers of an LLM:
Constrained Word Generation: Using standard prompts and the target word, we guide the model to generate the complete target word. At each layer, we track how often romanized tokens emerge during the decoding process. We do this by checking each generation timestep for latent romanization condition. The ‘latent fraction’ for a layer represents how frequently these romanized tokens appear across timesteps, averaged across all samples (c.f. Appendix C).
First Subword Only: We prompt for only the initial subword and compute the latent fraction. Despite having a single timestep, we maintain the latent fraction terminology for consistency.
Last Subword Only: We augment the standard prompt with all but the final subword of the target, then analyze the generation of the final subword.
We document layerwise latent fraction separately for first and last subword generation of the output. Intuitively, there is a distinction between the first and last token generation steps for a given word. In the former, the model faces a greater decision-making burden, while in the latter, the model is typically more confident in its predictions. We hypothesize that, in the latter scenario, the model may reach a decision in the layers just below the final few layers and express the output in a romanized form, as language-specific neurons, responsible for native script processing, are concentrated primarily in the last few layers Tang et al. (2024). This could lead to romanized tokens appearing more frequently as the model progresses from the first subword to the last.
4.2 Comparing Translations into Romanized vs. Native Script
This analysis examines translations into target languages in native script and their romanized equivalents, using the translation prompts from Section 4.1. We focus on first-token generation of the output word, also considering possible synonyms.
In the next token generation step, the probability of target language and latent language (English) Wendler et al. (2024) at each layer is examined using logit lens. Each probability is computed by summing over probabilities of all possible tokens corresponding to the answer word(s) in that respective language (c.f. Appendix D). Tokens of latent language and target language are derived using tokenization scheme for first token generation timestep mentioned in Section 4.1.
4.3 Activation Patching
Prompt design. The prompt template used here is identical to the translation prompt employed for the logit lens (Section 4.1).
Approach: We intend to compare how concepts are encoded in native script versus romanized script. In order to do this, we patch representations from source prompt where input language is romanized and compare this with patching from source prompt where input language is in native script. We do this for patching with single source prompt and repeat it for patching with mean over multiple source prompt, where multiple romanized source input languages are contrasted with multiple native script source input languages. Single-source patching might be influenced by language or script-specific characteristics. In contrast, multi-source patching reduces such biases, leading to more robust and generalizable findings.
In all scenarios from the resulting next token distribution, we compute the probabilities i.e probability of source concept in target language, and i.e probability of target concept in target language. We track , i.e., the probability of the concept C occurring in language , by simply summing up the probabilities of all prefixes of and its synonyms in the next-token distribution (c.f. Appendix E). We analyze to evaluate how effectively a concept is encoded in a given source input language .
5 Experimental Settings
5.1 Languages
We focus on five Indic languages: Hindi, Gujarati, Tamil, Telugu and Malayalam. Among these, Hindi and Gujarati belong to the Indo-Aryan branch of the Indo-European language family and use scripts derived from the Devanagari and Gujarati scripts, respectively. Tamil, Telugu, and Malayalam, on the other hand, are part of the Dravidian language family and each has its own distinct script. We also investigate latent romanization for Chinese and Georgian. Chinese belongs to the Sino-Tibetan language family and is written using logographic characters. Georgian is part of the Kartvelian language family and uses the unique Georgian script.
To examine the generality of latent romanization, we perform qualitative analyses on five additional languages that use different writing systems: Greek, Ukrainian, Amharic, Hebrew, and Arabic. Greek is part of the Hellenic branch of the Indo-European language family and is written using the Greek alphabet. Ukrainian belongs to the East Slavic group of the Indo-European family and employs the Ukrainian alphabet, a variant of the Cyrillic script. Amharic is a South Ethio-Semitic language within the Afroasiatic family and is written using the Ge’ez script, an abugida where each character represents a consonant-vowel combination. Hebrew is a Northwest Semitic language within the Afroasiatic family and is written using the Hebrew alphabet, an abjad script originating from the Aramaic alphabet. Arabic is a Central Semitic language, also part of the Afroasiatic family, and utilizes the Arabic script, another abjad that evolved from the Nabataean alphabet. Notably, both Hebrew and Arabic scripts are written from right to left.
5.2 Language Models
In this study, we focus mainly on Gemma-2 9b, Gemma-2 9b instruction-tuned Team et al. (2024), Llama-2 7b and Llama-2 13b Touvron et al. (2023) language models, some of the best performing open weights English-centric LLMs. Although the training data for these models are primarily English, these models have high multilingual capabilities Huang et al. (2023); Zhao et al. (2024a); Zhang et al. (2023).
| Native (input language in source prompt is in native script) | Romanized (input language in source prompt is in romanized script) | KL Divergence |
| Single Source | ||
| Multi Source | ||
5.3 Romanization
Transliteration is the conversion of text written in one script to another. Romanization is a subcategory of transliteration where the target script is English/Latin. We have taken into consideration two romanization schemes for Indic languages: (a) ITRANS scheme from IndicNLP library Kunchukuttan (2020) and (b) IndicXlit scheme Madhani et al. (2023). Based on our initial experiments, we observed that the IndicXlit scheme produces better romanization than ITRANS scheme. Hence for transliteration we have used the IndicXlit scheme Madhani et al. (2023). It generates romanization as is commonly used by native speakers and is trained on parallel transliteration corpora. We used pypinyin Mozillazg (2024) for Chinese and Unidecode Šolc (2025) for Georgian to romanize native scripts.
5.4 Logit Lens Experiments: Data
We use a curated word-level dataset with synonyms translated from recent works in this field Wendler et al. (2024) using the Llama 3.3 70B model Dubey et al. (2024). The quality of translations were manually verified to ensure accuracy and relevance. The datasets are kept simple to facilitate the observation of how latents evolve at each token level.
5.5 Activation Patching Experiments: Data
We adopt the dataset used in recent studies Dumas et al. (2024), extending it to include both native script and romanized versions of the languages considered in this study. Translations are performed using the Llama 3.3 70B model Dubey et al. (2024) and the translations are romanized using IndicXlit Madhani et al. (2023). All translations were manually validated to ensure data quality.
6 Results
6.1 Latent romanization
Our analysis demonstrates that LLMs do exhibit latent romanization during text completion tasks in six out of seven quantitatively analyzed languages, with Chinese being the only exception (c.f. 4) where this phenomenon is not observed. In Figure 4 we can see the latent fraction of romanized tokens for the last 10 layers of an LLM. This is across all tokens of the output word. The frequency of romanized tokens tends to increase just before the last layers. Qualitative logit lens analysis done for Greek, Ukrainian, Arabic, Amharic and Hebrew reveals similar patterns (c.f. Appendix F).
In Figure 5 it is observed that the range of the latent fraction of romanized tokens varies from in the first token generation step to in the last token generation step in most languages. This trend indicates that latent romanization increases progressively from the initial token to the final token of the output across languages. This observation supports our hypothesis that the first token generation involves more intricate decision-making processes compared to the generation of the final token within an output word.
Based on the above information, we quantify romanization across tasks for the last token generation step. Here the criteria would be if a romanized token occur at next token generation step with a probability > 0.1 in the last 10 layers, it is counted as a positive. As depicted in Figure 6, across translation, repetition and cloze task we observe a significant occurrence of romanized tokens in the latent layers.
Among the three text completion tasks considered in the latent romanization experiment (Section 4.1), a relatively high latent romanization frequency is observed for the repetition task (Figure 6). The translation and cloze tasks function at a semantic level, whereas the repetition task is purely syntactic. This means that the repetition task being less complex, the model may reach a decision sooner, potentially in earlier layers and might express its prediction in romanized form in intermediate layers. This behavior could be attributed to the distribution of language-specific neurons, responsible for native script processing, being predominantly concentrated in the initial and final few layers of LLMs Tang et al. (2024).
6.2 Patching with romanized representation versus native representation
As depicted in Figure 7, the probability distributions of source concepts in target language show remarkable similarity whether the source input language is in romanized or native script, consistent across both single-source and multi-source prompt setups. This analysis reveals that LLMs encode semantic concepts similarly regardless of whether the input is in native or romanized script. Furthermore, this finding demonstrates that the model achieves comparable levels of language understanding when processing non-Latin languages in their romanized form as in their native script.
6.3 Comparing Translations into Romanized vs. Native Script
We quantify the observations from Figure 2 by analyzing next-token predictions across layers using logit lens. In Figure 8, panels 8(b) and 8(d) illustrate that for translations into native scripts, target language tokens begin to emerge from layer 40 onward. Conversely, in panels 8(a) and 8(c), where the target language is in romanized script, target tokens appear 1–2 layers earlier. This pattern indicates that when processing non-Latin languages, the model forms internal representations of target tokens in earlier layers for romanized script compared to native script. This trend is consistent across language pairs and models, as shown in Figures 16-18. Additionally, native script curves (Figures 8(b), 8(d)) exhibit steeper gradients than their romanized equivalents (Figures 8(a), 8(c)).
7 Conclusion
In our investigation of romanized representations in the latent layers, we conclusively identified romanized tokens in the last 6–7 layers of an LLM across various multilingual text completion tasks. Based on previous works in this field Wendler et al. (2024); Zhao et al. (2024b), in an English-centric decoder only LLM this region corresponds to the transition from an English-centric language-agnostic concept space to a language-specific space where the idea conceived in the concept space is expressed in the target language. Our findings suggest that romanization serves as a bridge between the concept space and the language-specific region for non-Latin script languages, an observation strongly supported by our analysis of six diverse writing systems, though notably not observed in Chinese. Romanization acting as a bridge could explain why romanization based script barrier breaking methods like Liu et al. (2024); Xhelili et al. (2024) work.
Latent romanization is not observed for Chinese, unlike the other six languages examined in this study. This difference may be partially attributed to Chinese being a relatively high-resource language compared to the others analyzed, as well as its logographic script.
The identification of latent romanization in LLMs underscores its potential as a practical mechanism to address the challenges of limited training data for under-resourced languages. This perspective shifts the narrative from viewing such phenomena as an inherent English-centric bias to recognizing them as opportunities for innovation in multilingual AI. While our study provides initial insights into latent romanization, further research would be needed to understand whether and how these findings could inform training strategies for improving language model performance across diverse linguistic communities.
8 Limitations
The handling of multilingual text by large language models (LLMs) is an ongoing area of research. While evidence suggests that LLMs interact with English representations in a language-agnostic space, the mechanisms governing how this interaction varies at different timesteps during token generation of a word remain unclear. Our findings indicate that romanized representations become more prominent in hidden layers as the model progresses from the first to the last token generation step. This suggests that romanization may help the model address token fertility differences between output language and its primary latent language, English, especially for non-Latin languages with high token fertility. However, further research is needed to confirm and generalize these observations.
Furthermore, the scope of our analysis is restricted to models that utilize SentencePiece based tokenizers. The interpretability of non-Latin scripts at latent layers is limited when models employ tokenization schemes that split characters into multiple bytes, complicating logit lens analysis. Extending this work to models with alternative tokenization methods would offer a more complete understanding of multilingual capabilities and representations.
This work identifies but does not explain the selective occurrence and varying intensity of latent romanization across languages—questions that merit dedicated future investigation.
9 Ethics Statement
Through this work, our aim is to democratize access to LLMs and address the issue of limited data availability for low-resource languages. We emphasize that it is not our intention to diminish the value or significance of the native scripts of the languages included in this study.
Acknowledgments
We would like to thank EkStep Foundation and Nilekani Philanthropies for their generous grant towards research at AI4Bharat.
Appendix A English Translation and Romanization of the Sample Prompts
English translation and romanizations for the sample translation and repetition prompts mentioned in 4.1 are provided below.
English Translation.
Translation task.
Français: “poisson" Hindi: “fish Français: “mangue" Hindi: “mango Français: “frère" Hindi: “brother Français: “odeur" Hindi: “smell Français: “soleil" Hindi: “sun Français: “fleur" Hindi:
Repetition task.
Hindi: “fish Hindi: “fish Hindi: “mango Hindi: “mango Hindi: “brother Hindi: “brother Hindi: “smell Hindi: “smell Hindi: “sun Hindi: “sun Hindi: “flower Hindi:
Cloze task.
A “___" is used to play sports like soccer and basketball. Answer: “ball" A “___" is used for reading stories. Answer: “book" A “___" is often given as a gift and can be found in gardens. Answer:
English Transliteration.
Translation task.
Français: “poisson" Hindi: “machhalee Français: “mangue" Hindi: “aam Français: “frère" Hindi: “bhaee Français: “odeur" Hindi: “gandh Français: “soleil" Hindi: “sooraj Français: “fleur" Hindi:
Repetition task.
Hindi: “machhalee Hindi: “machhalee Hindi: “aam Hindi: “aam Hindi: “bhaee Hindi: “bhaee Hindi: “gandh Hindi: “gandh Hindi: “sooraj Hindi: “sooraj Hindi: “phool Hindi:
Cloze task.
Phutball aur baasketball jaise khel khelane ke lie “___" ka upayog kiya jaata hai. Uttar: “ball" Ek “___" kahaaniyaan padhane ke lie upayog kiya jaata hai. Uttar: “ball" Ek “___" aksar upahaar ke roop mein diya jaata hai aur yah bageechon mein paaya ja sakata hai. Uttar:
Appendix B Latent romanization: Tokenization scheme for the romanized word
Mathematically, we track the following romanized tokens for a given romanized word of length :
First timestep: , "_ " represents single leading space and represents prefixes of Intermediate timesteps: , represents sub-strings of Final timestep: , , represents suffixes of
Similarly we construct token sets for the native script () and English () by including prefixes of the corresponding word, both with and without leading spaces, for all timesteps except the last. For the final timestep, we use suffixes of the corresponding word. We discard the sample if there is any overlap between romanized tokens () and either the native () or English tokens () as shown by the following condition:
[…]
Appendix C Latent Fraction
Formally, we compute the latent fraction as follows:
For layer , timestep , sample and set of corresponding romanized tokens : 1. Latent romanization condition: 2. Latent fraction for a layer : where is the number of samples, is the number of generation timesteps and is the probability of generating token at timestep and layer .
Appendix D Computing Language probabilities - For translation towards native script vs translation towards romanized script task
To compute language probabilities, we search the LLM’s vocabulary for all tokens that could be the first token of the correct ouput word(s) in the respective language. We search the models vocabulary for all prefixes of the word(s) with and without leading space. For a language with corresponding output word and its synonyms , we define:
where is the set of all prefixes of output word and its synonyms , including versions with and without leading spaces. For example to get probability for english when the output word is “fast and its synonym is “swift, then (all the token-level prefixes of “fast, “_fast, “swift and “_swift). “_ represents a single leading space.
Appendix E Computing Probabilities : Activation Patching Experiment
Probability for a concept in language can be formulated as :
where is the set of all prefixes of output word and its synonyms (note that a word’s tokens are its prefixes).
We keep source concept and target concept distinct to avoid ambiguity when both are expressed in the same target language .
Cases of token overlap between i.e word representing source concept in target language and i.e word representing target concept in target language and their synonyms are excluded. Token overlap would cause ambiguity. Therefore in the final dataset,
Where represents all the prefixes of and its synonyms.
Appendix F Latent Romanization Qualitative Analysis
We list qualitative logit lens analysis for Greek, Ukrainian, Hebrew, Arabic and Amharic (see Figures 9 to 13).
, ‘Bab’ is the romanized form ) taking shape using logit lens producing a next-token distribution for each position (x-axis) and layers 20 and above (y-axis).Interestingly in middle to top layers (20 - 29) we could observe romanized subwords of the Arabic word (
- Bab;
- a ;
- ab) before they are represented in their native script. Color represents entropy of next-token generation from low (blue) to high (red). Plotting tool: Belrose et al. (2023).
, ‘muzika’ is the romanized form ) taking shape using logit lens producing a next-token distribution for each position (x-axis) and layers 24 and above (y-axis).Interestingly in middle to top layers we could observe romanized subwords of the Amharic word (
- muzy;
- z) before they are represented in their native script. Color represents entropy of next-token generation from low (blue) to high (red). Plotting tool: Belrose et al. (2023).
Appendix G Latent Romanization Quantitative Analysis
Quantitative Analysis of latent romanization for repetition task with gemma 2 9b it model can be seen in Figure 14. Layerwise fractional distribution of romanized tokens across output token generation timesteps for translation and repetition task with Gemma 2 9b, Llama 2 7b, and Llama 2 13b models are present in Figure 15.
| Gemma 2 9b | Llama 2 13b | Llama 2 7b |
Translation
Repetition
Appendix H Translation towards native script vs Translation towards romanized script: Additional examples
Translation towards native script is compared with translation towards romanized script for gemma 2 9b it, gemma 2 9b and llama 2 13b models (see Figures 16 to 18).

### v2

@@
-RomanLens: Latent Romanization and its role in Multilinguality in LLMs
+RomanLens: The Role Of Latent Romanization In Multilinguality In LLMs
 Abstract
@@
 A fundamental question arises: how do LLMs achieve such robust multilingual capabilities?
-For non-Latin script languages, we investigate the role of romanization—the representation of non-Latin scripts using Latin characters—as a bridge in multilingual processing.
+We take the case of non-Roman script languages, we investigate the role of Romanization—the representation of non-Roman scripts using Roman characters—as a bridge in multilingual processing.
 Using mechanistic interpretability techniques, we analyze next-token generation and find that intermediate layers frequently represent target words in Romanized form before transitioning to native script, a phenomenon we term Latent Romanization.
 Further, through activation patching experiments, we demonstrate that LLMs encode semantic concepts similarly across native and Romanized scripts, suggesting a shared underlying representation.
-Additionally in translation towards non Latin languages, our findings reveal that when the target language is in romanized form, its representations emerge earlier in the model’s layers compared to native script.
+Additionally, for translation into non-Roman script languages, our findings reveal that when the target language is in Romanized form, its representations emerge earlier in the model’s layers compared to native script.
 These insights contribute to a deeper understanding of multilingual representation in LLMs and highlight the implicit role of Romanization in facilitating language transfer.
-Our work provides new directions for potentially improving multilingual language modeling and interpretability.
-RomanLens: Latent Romanization and its role in Multilinguality in LLMs
+RomanLens: The Role Of Latent Romanization In Multilinguality In LLMs
 Alan Saji1222Work done during employment at AI4Bharat.333
@@
 Raj Dabre1,3,4,5
-Anoop Kunchukuttan1,3,6,
-Mitesh M. Khapra1,3,
+Anoop Kunchukuttan1,6,
 Ratish Puduppully7333
@@
 (2024) describe distinct phases in multilingual information processing: initial layers map language-specific lexical and syntactic representations to a language-independent semantic space, middle layers maintain this semantic abstraction, and final layers transform these representations into language-specific lexical and syntactic forms.
-In this study, we focus on languages written in non-Latin scripts.
-We hypothesize that LLMs leverage romanized forms of non-Latin script languages as an intermediate bridge between their language-agnostic concept space and language-specific output representations.
-Romanization—the representation of non-Latin scripts using Latin characters—may facilitate this process by aligning non-English languages more closely with English.
-Supporting this, Husain et al.
+In this study, we focus on languages written in non-Roman scripts.
+We hypothesize that LLMs leverage romanized forms of non-Roman script languages as an intermediate bridge between their language-agnostic concept space and language-specific output representations.
+Romanization—the representation of non-Roman scripts using Roman characters—may facilitate this process by aligning non-English languages more closely with English.
+Supporting this, Jaavid et al.
 (2024) demonstrated that explicitly romanizing inputs improves model performance on multilingual tasks, suggesting an inherent alignment between romanized text and English representations.
@@
 (2024); Dumas et al.
-(2024), a technique that replaces activations from one forward pass with another to analyze the resulting outputs (see Section 3.2).
+(2024), a technique that replaces activations from one forward pass with another to analyze the resulting outputs (c.f.
+Section 3.2).
 Dumas et al.
@@
 Building on this, we perform layerwise activation patching between romanized and native-script inputs to examine whether LLMs encode conceptual information similarly across scripts.
-, ’pa’ is its romanized representation). Note that push is a prefix of pushpam, the romanized form of the translation of “flower”; also, poo is another romanized translation of “flower.”
+, ’pa’ is its romanized representation).
+Note that push is a prefix of pushpam, the romanized form of the translation of “flower”; also, poo is another romanized translation of “flower”.Color represents entropy of next-token generation from low (blue) to high (red).
+Plotting tool: Belrose et al.
+(2023).
 Based on our experiments, we list down a summary of our contributions below:
@@
 We term this phenomenon Latent Romanization.
-2.
-Consistent Semantic Encoding Across Scripts: Activation patching experiments reveal that LLMs encode semantic concepts similarly, regardless of whether the input is in native or romanized script.
+2. Consistent Semantic Encoding Across Scripts: Activation patching experiments reveal that LLMs encode semantic concepts similarly, regardless of whether the input is in native or Romanized script.
 3.
-Earlier Emergence of Target Representations: When translating into romanized versus native script, romanized target representations emerge earlier in the model’s layers—typically one or two layers prior to native script representations (Figure 2).
-This suggests that romanization facilitates faster progression toward language-specific embeddings.
+Earlier Emergence of Target Representations: When translating into Romanized versus native script, Romanized target representations emerge earlier in the model’s layers—typically one or two layers prior to native script representations (c.f.
+Figure 2).
 2 Related Work
-Recent studies investigate latent representations in multilingual tasks, examining whether English emerges as a latent language in English-centric LLMs Wendler et al.
-(2024) and how the composition of training corpus mixtures influences latent representations Zhong et al.
-(2024).
-Recent investigations analyze how LLMs handle multilingual capabilitiesZhao et al.
+Recent studies have explored various aspects of LLMs’ multilingual behavior: examining whether English emerges as a latent language in English-centric LLMs Wendler et al.
+(2024), how the composition of training corpus mixtures influences latent representations Zhong et al.
+(2024) and how LLMs handle multilingual capabilitiesZhao et al.
 (2024b).
@@
 Direct logit attribution attributes logits to individual attention heads.
-This work focuses on the logit lens (Section 4.1) to investigate whether English-centric decoder only LLMs like Gemma-2, when prompted in non-latin language, processes via romanized latent states before producing non-English text.
-Tuned lens is avoided as its training process might obscure the intermediate romanized states by aligning them to final non-English outputs, potentially masking the phenomenon under investigation.
+This work focuses on the logit lens (Section 4.1) to investigate whether English-centric decoder only LLMs when prompted in a non-Roman language, processes via romanized latent states before producing native language text.
+Tuned lens is avoided as its training process might obscure the intermediate romanized states by aligning them to final native script outputs, potentially masking the phenomenon under investigation.
 Activation patching (Meng et al., 2022) is a key interpretability technique employed in our study.
@@
 (2024).
-Building on these approaches, we adopt an activation patching-based experimental framework as defined in Dumas et al.
-(2024) to investigate and compare how concepts are encoded in romanized versus native scripts.
-Previous studies have demonstrated that romanization can serve as an effective approach to interact with LLMs Husain et al.
+Building on these approaches, we adopt an activation patching-based experimental framework to investigate and compare how concepts are encoded in romanized versus native scripts.
+Previous studies have demonstrated that romanization can serve as an effective approach to interact with LLMs Jaavid et al.
 (2024).
@@
 (2024) and Xhelili et al.
-(2024) employ an approach based on contrastive learning for post-training alignment, contrasting sentences with their transliterations in Latin script to overcome the script barrier and enhance cross-lingual transfer.
+(2024) employ an approach based on contrastive learning for post-training alignment, contrasting sentences with their transliterations in Roman script to overcome the script barrier and enhance cross-lingual transfer.
 However, our work distinguishes itself from prior research by exploring the presence of romanized representations in the latent layers of an LLM during multilingual tasks, an aspect that, to the best of our knowledge, has not yet been investigated.
@@
 We give a quick background of the transformer’s forward pass as well as the basics of mechanistic interpretability approaches such as logit lens and activation patching which we leverage in this paper.
-Transformer’s forward pass: Here, we study autoregressive decoder-only transformer models.
-Vaswani (2017).
-Such models maintain the dimensionality of the input data throughout the computation process.
-For an input sequence , where is the sequence length, each token is initially mapped to an embedding vector, which is then transformed through the model’s layers, producing a sequence of hidden states (also referred to as latents).
-The initial latents are obtained from a learned embedding matrix, which maps each vocabulary token to a fixed vector.
-These latents are updated iteratively across layers through the addition of residuals.
-The update rule for the latent at position in layer is expressed as:
-The function , , is known as a transformer block, where is the final layer index.
-For prediction, the final embedding vector is multiplied by an unembedding matrix , producing a vector containing the logit scores for each vocabulary token .
-This is computed as:
-These logit scores are converted to probabilities via the softmax function as follows:
+Transformer’s Forward Pass
+Decoder-only transformer models (Vaswani, 2017) employ a residual architecture to process input sequences through multiple layers, producing a sequence of hidden states (latents).
+These latents, whose dimensionality remains the same are updated iteratively across layers through transformer blocks , where indicates the layer index and is the final layer index.
+For next-token prediction, the final latent is transformed by an unembedding matrix to produce logit scores for vocabulary tokens which are then converted to probabilities via the softmax function (c.f.
+Appendix A).
 3.1 Interpretability Tool: Logit lens
@@
 We design our analysis setup with the intention of addressing the following research questions:
-RQ1: Does LLMs exhibit latent romanization during multilingual text completion tasks? (Section 4.1)
-RQ2: What are the differences in hidden layer representations when processing the same language in romanized and native scripts? (Section 4.2)
-RQ3: How does the representation of semantic concepts in LLMs compare between native and romanized scripts of non-Latin languages? (Section 4.3)
-4.1 Logit lens
+RQ1: Do LLMs exhibit latent romanization during multilingual text completion tasks? (Section 4.1)
+RQ2: How does the representation of semantic concepts in LLMs compare between native and romanized scripts of non-Roman languages? (Section 4.2)
+RQ3: What are the differences in hidden layer representations when processing the same language in romanized and native scripts? (Section 4.3)
 Prompt design.
-In order to have the next token be easily inferred from the given prompt , we establish the following prompts.We use Hindi as the example language to describe the prompts.
-English translation and transliteration of the below prompts are present in Appendix A.
-[…]
-[…]
-Repetition task. We ask the model to repeat the word mentioned in the same language.
-[…]
-Cloze task. We ask the model to predict the masked word in a sentence. We give the model two examples and ask it to predict the masked word in the next masked sentence.
-[…]
-Latent Romanization Analysis
+We design prompts that facilitate next-token () prediction from the given context ().
+This is adopted across all analysis setups.
+Appendix B provides a Hindi example of this prompt design, including its English translation and romanized form.
+Translation task. We prompt the model to translate a word given five in-context examples.
+Repetition task. We prompt the model to repeat a word in the same language given 5 in-context examples.
+Cloze task. We prompt the model to predict the masked word in a sentence given two in-context examples.
+4.1 Latent Romanization Analysis
 Translation, repetition, and cloze tasks are explored by providing the respective prompts as inputs to an LLM to generate the corresponding output word.
 We romanize the output word, tokenize it, retaining only the tokens present in the model’s vocabulary, and analyze the occurrence of these tokens in the latent layers across timesteps of the output word generation.
-The analysis is done using logit lens by examining whether the probability of a romanized token in the next token distribution at a given layer exceeds 0.1 - a threshold optimized for detection accuracy i.e minimizing false positives and maximizing true positives.
+The analysis is done using logit lens by examining whether the probability of a romanized token in the next token distribution at a given layer exceeds 0.1.
 We refer to this hereafter as the latent romanization condition.
+The 0.1 threshold is empirically determined to optimize detection accuracy, i.e.
+minimizing false positives and maximizing true positives (compared to alternative thresholds 0.05 and 0.01).
 Our analysis focuses on the final 10 layers of an LLM, where coherent romanized representations emerge according to logit lens visualizations (c.f.
-Figures 1 and 9 - 13).
+Figures 1 and 8 - 12).
 We track romanized tokens using a timestep-specific tokenization scheme optimized for detection accuracy.
@@
 In the final output generation timestep, we probe the presence of only the full romanized word and its suffixes as potential tokens (c.f.
-Appendix B).
-Latent romanization is analyzed under three distinct scenarios for the last 10 layers of an LLM:
+Appendix C).
+Latent romanization is analyzed under three distinct scenarios:
 Constrained Word Generation: Using standard prompts and the target word, we guide the model to generate the complete target word.
@@
 The ‘latent fraction’ for a layer represents how frequently these romanized tokens appear across timesteps, averaged across all samples (c.f.
-Appendix C).
+Appendix D).
 First Subword Only: We prompt for only the initial subword and compute the latent fraction. Despite having a single timestep, we maintain the latent fraction terminology for consistency.
@@
 This could lead to romanized tokens appearing more frequently as the model progresses from the first subword to the last.
+4.2 Patching With Romanized Representation Versus Native Representation
+We intend to compare how concepts are encoded in native script versus romanized script using translation task.
+In order to do this, we patch representations from source prompt where input language is romanized and compare this with patching from source prompt where input language is in native script.
+We first perform patching using a single-source prompt.
+Then, we repeat the process using an averaged multi-source prompt, contrasting multiple romanized source input languages with multiple native script source input languages.
+Single-source patching might be influenced by language or script-specific characteristics.
+In contrast, multi-source patching reduces such biases, leading to more robust and generalizable findings.
+In all scenarios from the resulting next token distribution, we compute the probabilities i.e probability of source concept in target language, and i.e probability of target concept in target language.
+We track , i.e., the probability of the concept C occurring in language , by simply summing up the probabilities of all prefixes of and its synonyms in the next-token distribution (c.f.
+Appendix G).
+We analyze to evaluate how effectively a concept is encoded in a given source input language .
 4.3 Comparing Translations Into Romanized vs. Native Script
-This analysis examines translations into target languages in native script and their romanized equivalents, using the translation prompts from Section 4.1.
+This analysis examines translation task with target languages in native script and their romanized equivalents.
 We focus on first-token generation of the output word, also considering possible synonyms.
@@
 Each probability is computed by summing over probabilities of all possible tokens corresponding to the answer word(s) in that respective language (c.f.
-Appendix D).
+Appendix E).
 Tokens of latent language and target language are derived using tokenization scheme for first token generation timestep mentioned in Section 4.1.
-4.3 Activation Patching
-Prompt design. The prompt template used here is identical to the translation prompt employed for the logit lens (Section 4.1).
-Approach: We intend to compare how concepts are encoded in native script versus romanized script.
-In order to do this, we patch representations from source prompt where input language is romanized and compare this with patching from source prompt where input language is in native script.
-We do this for patching with single source prompt and repeat it for patching with mean over multiple source prompt, where multiple romanized source input languages are contrasted with multiple native script source input languages.
-Single-source patching might be influenced by language or script-specific characteristics.
-In contrast, multi-source patching reduces such biases, leading to more robust and generalizable findings.
-In all scenarios from the resulting next token distribution, we compute the probabilities i.e probability of source concept in target language, and i.e probability of target concept in target language.
-We track , i.e., the probability of the concept C occurring in language , by simply summing up the probabilities of all prefixes of and its synonyms in the next-token distribution (c.f.
-Appendix E).
-We analyze to evaluate how effectively a concept is encoded in a given source input language .
 5 Experimental Settings
-5.1 Languages
-We focus on five Indic languages: Hindi, Gujarati, Tamil, Telugu and Malayalam.
+Languages:
+We focus on five Indic languages: Hindi, Gujarati, Tamil, Telugu, and Malayalam, as well as Chinese and Georgian.
 Among these, Hindi and Gujarati belong to the Indo-Aryan branch of the Indo-European language family and use scripts derived from the Devanagari and Gujarati scripts, respectively.
 Tamil, Telugu, and Malayalam, on the other hand, are part of the Dravidian language family and each has its own distinct script.
-We also investigate latent romanization for Chinese and Georgian.
 Chinese belongs to the Sino-Tibetan language family and is written using logographic characters.
 Georgian is part of the Kartvelian language family and uses the unique Georgian script.
-To examine the generality of latent romanization, we perform qualitative analyses on five additional languages that use different writing systems: Greek, Ukrainian, Amharic, Hebrew, and Arabic.
-Greek is part of the Hellenic branch of the Indo-European language family and is written using the Greek alphabet.
-Ukrainian belongs to the East Slavic group of the Indo-European family and employs the Ukrainian alphabet, a variant of the Cyrillic script.
-Amharic is a South Ethio-Semitic language within the Afroasiatic family and is written using the Ge’ez script, an abugida where each character represents a consonant-vowel combination.
-Hebrew is a Northwest Semitic language within the Afroasiatic family and is written using the Hebrew alphabet, an abjad script originating from the Aramaic alphabet.
-Arabic is a Central Semitic language, also part of the Afroasiatic family, and utilizes the Arabic script, another abjad that evolved from the Nabataean alphabet.
-Notably, both Hebrew and Arabic scripts are written from right to left.
-5.2 Language Models
+To examine the generality of latent romanization, we perform qualitative analyses on five additional languages that use different writing systems: Greek, Ukrainian, Amharic, Hebrew, and Arabic (c.f Appendix H).
+Language Models:
 In this study, we focus mainly on Gemma-2 9b, Gemma-2 9b instruction-tuned Team et al.
-(2024), Llama-2 7b and Llama-2 13b Touvron et al.
-(2023) language models, some of the best performing open weights English-centric LLMs.
+(2024), Llama-2 7b, Llama-2 13b Touvron et al.
+(2023) and Mistral- 7B Jiang et al.
+(2023) (c.f.
+Appendix K) language models, some of the best performing open weights English-centric LLMs.
 Although the training data for these models are primarily English, these models have high multilingual capabilities Huang et al.
@@
 (2023).
-| Native (input language in source prompt is in native script) | Romanized (input language in source prompt is in romanized script) | KL Divergence |
-| Single Source | ||
-| Multi Source | ||
-5.3 Romanization
-Transliteration is the conversion of text written in one script to another.
-Romanization is a subcategory of transliteration where the target script is English/Latin.
-We have taken into consideration two romanization schemes for Indic languages: (a) ITRANS scheme from IndicNLP library Kunchukuttan (2020) and (b) IndicXlit scheme Madhani et al.
-(2023).
-Based on our initial experiments, we observed that the IndicXlit scheme produces better romanization than ITRANS scheme.
-Hence for transliteration we have used the IndicXlit scheme Madhani et al.
-(2023).
-It generates romanization as is commonly used by native speakers and is trained on parallel transliteration corpora.
-We used pypinyin Mozillazg (2024) for Chinese and Unidecode Šolc (2025) for Georgian to romanize native scripts.
-5.4 Logit Lens Experiments: Data
+Romanization:
+Data For Logit Lens Experiments:
 We use a curated word-level dataset with synonyms translated from recent works in this field Wendler et al.
@@
 The datasets are kept simple to facilitate the observation of how latents evolve at each token level.
-5.5 Activation Patching Experiments: Data
+Data For Activation Patching Experiments:
 We adopt the dataset used in recent studies Dumas et al.
@@
 6 Results
+We now share our observations about the role of Romanization in LLMs.
+| Native | Romanized | |
+| Single Source | ||
+| Multi Source | ||
 6.1 Latent romanization
-Our analysis demonstrates that LLMs do exhibit latent romanization during text completion tasks in six out of seven quantitatively analyzed languages, with Chinese being the only exception (c.f.
-4) where this phenomenon is not observed.
-In Figure 4 we can see the latent fraction of romanized tokens for the last 10 layers of an LLM.
+Our analysis demonstrates that LLMs do exhibit latent romanization during text completion tasks in six out of seven quantitatively analyzed languages (c.f.
+Figure 4(a)), with Chinese being the only exception where this phenomenon is not observed.
+In Figure 4(a) we can see the latent fraction of romanized tokens for the last 10 layers of an LLM.
 This is across all tokens of the output word.
@@
 Qualitative logit lens analysis done for Greek, Ukrainian, Arabic, Amharic and Hebrew reveals similar patterns (c.f.
-Appendix F).
-In Figure 5 it is observed that the range of the latent fraction of romanized tokens varies from in the first token generation step to in the last token generation step in most languages.
+Appendix H).
+In Figures 4(b) and 4(c) it is observed that the range of the latent fraction of romanized tokens varies from in the first token generation step to in the last token generation step in most languages.
 This trend indicates that latent romanization increases progressively from the initial token to the final token of the output across languages.
@@
 This means that the repetition task being less complex, the model may reach a decision sooner, potentially in earlier layers and might express its prediction in romanized form in intermediate layers.
-This behavior could be attributed to the distribution of language-specific neurons, responsible for native script processing, being predominantly concentrated in the initial and final few layers of LLMs Tang et al.
+This behavior could be attributed to language-specific neurons, responsible for native script processing, being predominantly concentrated in the initial and final few layers of LLMs Tang et al.
 (2024).
-6.2 Patching with romanized representation versus native representation
-As depicted in Figure 7, the probability distributions of source concepts in target language show remarkable similarity whether the source input language is in romanized or native script, consistent across both single-source and multi-source prompt setups.
+6.2 Patching With Romanized Representation vs. Native Representation
+In Figure 5, we analyze two patching scenarios: In the single-source setup, we compare patching from Hindi→Italian with Hindi(romanized)→Italian source prompt, to Malayalam→Italian target prompt.
+In the multi-source setup, we contrast patching from multiple native script prompts (Hindi→Italian, Gujarati→Italian ) against their romanized counterparts (Hindi(romanized)→Italian, Gujarati(romanized)→Italian ).
+We compare the probability distributions of source concept in target language across adjacent graphs where native source input language is contrasted with romanized source input language.
+It is evident that these probability distributions show remarkable similarity whether the source input language is in romanized or native script, consistent across both single-source and multi-source prompt setups.
+The similarity is quantitatively supported by the KL divergence measurements between adjacent graphs, remaining below 0.01 in both setups.
+KL divergence value close to zero indicates that the two distributions are nearly identical.
 This analysis reveals that LLMs encode semantic concepts similarly regardless of whether the input is in native or romanized script.
-Furthermore, this finding demonstrates that the model achieves comparable levels of language understanding when processing non-Latin languages in their romanized form as in their native script.
+Furthermore, this finding demonstrates that the model achieves comparable levels of language understanding when processing non-Roman languages in their romanized form as in their native script.
 6.3 Comparing Translations Into Romanized vs. Native Script
 We quantify the observations from Figure 2 by analyzing next-token predictions across layers using logit lens.
-In Figure 8, panels 8(b) and 8(d) illustrate that for translations into native scripts, target language tokens begin to emerge from layer 40 onward.
-Conversely, in panels 8(a) and 8(c), where the target language is in romanized script, target tokens appear 1–2 layers earlier.
-This pattern indicates that when processing non-Latin languages, the model forms internal representations of target tokens in earlier layers for romanized script compared to native script.
+In Figure 7, panels 7(c) and 7(d) illustrate that for translations into native scripts, target language tokens begin to emerge from layer 40 onward.
+Conversely, in panels 7(a) and 7(b), where the target language is in romanized script, target tokens appear 1–2 layers earlier.
+This pattern indicates that when processing non-Roman script languages, the model forms internal representations of target tokens in earlier layers for romanized script compared to native script.
 This trend is consistent across language pairs and models, as shown in Figures 16-18.
-Additionally, native script curves (Figures 8(b), 8(d)) exhibit steeper gradients than their romanized equivalents (Figures 8(a), 8(c)).
-7 Conclusion
+This suggests that romanization facilitates faster progression toward language-specific embeddings.
+In Figure 7, in all four graphs consistent with Wendler et al.
+(2024), English representations emerge from the middle layers and persist until the final few layers, where the target language representations gradually take shape.
+It is important to note that native script curves (Figures 7(c), 7(d)) exhibit steeper gradients than their romanized equivalents (Figures 7(a), 7(b)).
+Discussion.
 In our investigation of romanized representations in the latent layers, we conclusively identified romanized tokens in the last 6–7 layers of an LLM across various multilingual text completion tasks.
@@
 (2024b), in an English-centric decoder only LLM this region corresponds to the transition from an English-centric language-agnostic concept space to a language-specific space where the idea conceived in the concept space is expressed in the target language.
-Our findings suggest that romanization serves as a bridge between the concept space and the language-specific region for non-Latin script languages, an observation strongly supported by our analysis of six diverse writing systems, though notably not observed in Chinese.
+Our findings suggest that romanization serves as a bridge between the concept space and the language-specific region for non-Roman script languages, an observation strongly supported by our analysis of six diverse writing systems.
 Romanization acting as a bridge could explain why romanization based script barrier breaking methods like Liu et al.
@@
 (2024) work.
-Latent romanization is not observed for Chinese, unlike the other six languages examined in this study.
-This difference may be partially attributed to Chinese being a relatively high-resource language compared to the others analyzed, as well as its logographic script.
-The identification of latent romanization in LLMs underscores its potential as a practical mechanism to address the challenges of limited training data for under-resourced languages.
-This perspective shifts the narrative from viewing such phenomena as an inherent English-centric bias to recognizing them as opportunities for innovation in multilingual AI.
-While our study provides initial insights into latent romanization, further research would be needed to understand whether and how these findings could inform training strategies for improving language model performance across diverse linguistic communities.
+Notably, we do not observe Latent Romanization in Chinese, likely due to its logographic script and relatively high-resource status.
+7 Conclusion
+Our findings show that LLMs implicitly use Romanization as a bridge for non-Roman scripts, exhibiting Latent Romanization in intermediate layers before switching to native scripts.
+Layerwise analyses reveal that semantic concepts are encoded similarly across native and Romanized inputs, indicating a shared internal representation.
+Moreover, when translating to a Romanized script, target words emerge earlier, highlighting Romanization as a structural link between language-agnostic concepts and language-specific output.
+While our study reveals initial insights into Latent Romanization, future work could focus on applying these findings to develop training strategies that enhance performance across diverse linguistic communities.
 8 Limitations
-The handling of multilingual text by large language models (LLMs) is an ongoing area of research.
-While evidence suggests that LLMs interact with English representations in a language-agnostic space, the mechanisms governing how this interaction varies at different timesteps during token generation of a word remain unclear.
-Our findings indicate that romanized representations become more prominent in hidden layers as the model progresses from the first to the last token generation step.
-This suggests that romanization may help the model address token fertility differences between output language and its primary latent language, English, especially for non-Latin languages with high token fertility.
+The handling of multilingual text by large language models (LLMs) remains an active area of research.
+Although evidence suggests that LLMs process English representations within a language-agnostic space, the specific mechanisms by which these models adjust their interactions over different timesteps during token generation are still not fully understood.
+In our study, we observe that romanized representations become increasingly prominent in the hidden layers as token generation progresses from the first to the final token.
+This trend suggests that latent romanization may help the model mitigate differences in token fertility—that is, the average number of tokens required to represent a word—between the output language and its primary latent language, English.
+This effect appears especially for non-Roman script languages, with high token fertility.
 However, further research is needed to confirm and generalize these observations.
-Furthermore, the scope of our analysis is restricted to models that utilize SentencePiece based tokenizers.
-The interpretability of non-Latin scripts at latent layers is limited when models employ tokenization schemes that split characters into multiple bytes, complicating logit lens analysis.
+The interpretability of non-Roman scripts at latent layers is limited when models employ tokenization schemes that split non-roman characters into multiple bytes, complicating logit lens analysis.
 Extending this work to models with alternative tokenization methods would offer a more complete understanding of multilingual capabilities and representations.
@@
 We emphasize that it is not our intention to diminish the value or significance of the native scripts of the languages included in this study.
+The code and datasets created in this work will be made available under permissible licenses.
+Generative AI systems were only used for assistance purely with the language of the paper, eg: paraphrasing, spell-check, polishing the author’s original content, and for writing boiler-plate code.
 Acknowledgments
-We would like to thank EkStep Foundation and Nilekani Philanthropies for their generous grant towards research at AI4Bharat.
-Appendix A English Translation and Romanization of the Sample Prompts
-English translation and romanizations for the sample translation and repetition prompts mentioned in 4.1 are provided below.
+Appendix A Transformer’s Forward pass: Detailed
+For an input sequence , where is the sequence length, the initial latents are obtained from a learned embedding matrix. The update rule for the latent at position in layer is expressed as:
+The logit scores are computed as:
+These are converted to probabilities via the softmax function:
+Appendix B Sample Prompts
+A Hindi example, its English translation and transliteration for the translation, repetition and cloze task prompt designs mentioned in Section 4 are provided below.
+Hindi example.
+Translation task. A translation prompt from French to Hindi.
+[…]
+Repetition task.
+[…]
+Cloze task..
+[…]
 English Translation.
@@
 Uttar:
-Appendix B Latent romanization: Tokenization scheme for the romanized word
+Appendix C Latent romanization: Tokenization scheme for the romanized word
 Mathematically, we track the following romanized tokens for a given romanized word of length :
@@
 […]
-Appendix C Latent Fraction
+Appendix D Latent Fraction
 Formally, we compute the latent fraction as follows:
@@
 Latent fraction for a layer : where is the number of samples, is the number of generation timesteps and is the probability of generating token at timestep and layer .
-Appendix D Computing Language probabilities - For translation towards native script vs translation towards romanized script task
+Appendix E Computing Language probabilities - For translation towards native script vs translation towards romanized script task
 To compute language probabilities, we search the LLM’s vocabulary for all tokens that could be the first token of the correct ouput word(s) in the respective language.
@@
 “_ represents a single leading space.
-Appendix E Computing Probabilities : Activation Patching Experiment
+Appendix F Romanization scheme: Indic languages
+We have taken into consideration two romanization schemes for Indic languages: (a) ITRANS scheme from IndicNLP library Kunchukuttan (2020) and (b) IndicXlit scheme Madhani et al.
+(2023).
+Based on our initial experiments, we observed that the IndicXlit scheme produces better romanization than ITRANS scheme.
+Thus for romanization we have used the IndicXlit scheme Madhani et al.
+(2023).
+It generates romanization as is commonly used by native speakers and is trained on parallel transliteration corpora.
+Appendix G Computing Probabilities : Activation Patching Experiment
 Probability for a concept in language can be formulated as :
@@
 Where represents all the prefixes of and its synonyms.
-Appendix F Latent Romanization Qualitative Analysis
-We list qualitative logit lens analysis for Greek, Ukrainian, Hebrew, Arabic and Amharic (see Figures 9 to 13).
+Appendix H Latent Romanization Qualitative Analysis
+We list qualitative logit lens analysis for Greek, Ukrainian, Hebrew, Arabic and Amharic (see Figures 8 to 12).
+Languages.
+Greek is part of the Hellenic branch of the Indo-European language family and is written using the Greek alphabet.
+Ukrainian belongs to the East Slavic group of the Indo-European family and employs the Ukrainian alphabet, a variant of the Cyrillic script.
+Amharic is a South Ethio-Semitic language within the Afroasiatic family and is written using the Ge’ez script, an abugida where each character represents a consonant-vowel combination.
+Hebrew is a Northwest Semitic language within the Afroasiatic family and is written using the Hebrew alphabet, an abjad script originating from the Aramaic alphabet.
+Arabic is a Central Semitic language, also part of the Afroasiatic family, and utilizes the Arabic script, another abjad that evolved from the Nabataean alphabet.
+Notably, both Hebrew and Arabic scripts are written from right to left.
 , ‘Bab’ is the romanized form ) taking shape using logit lens producing a next-token distribution for each position (x-axis) and layers 20 and above (y-axis).Interestingly in middle to top layers (20 - 29) we could observe romanized subwords of the Arabic word (
@@
 - z) before they are represented in their native script. Color represents entropy of next-token generation from low (blue) to high (red). Plotting tool: Belrose et al. (2023).
-Appendix G Latent Romanization Quantitative Analysis
-Quantitative Analysis of latent romanization for repetition task with gemma 2 9b it model can be seen in Figure 14.
-Layerwise fractional distribution of romanized tokens across output token generation timesteps for translation and repetition task with Gemma 2 9b, Llama 2 7b, and Llama 2 13b models are present in Figure 15.
+Appendix I Latent Romanization Quantitative Analysis: Additional examples
+Quantitative Analysis of latent romanization for repetition task and cloze task with gemma 2 9b it model can be seen in Figures 13 and 14 respectively.
+Layerwise fractional distribution of romanized tokens across output token generation timesteps for translation, repetition and cloze task with Gemma 2 9b, Llama 2 7b, and Llama 2 13b models are present in Figure 15.
 | Gemma 2 9b | Llama 2 13b | Llama 2 7b |
@@
 Repetition
-Appendix H Translation towards native script vs Translation towards romanized script: Additional examples
+Cloze
+Appendix J Comparing Translations Into Romanized vs. Native Script: Additional examples
 Translation towards native script is compared with translation towards romanized script for gemma 2 9b it, gemma 2 9b and llama 2 13b models (see Figures 16 to 18).
+Appendix K Other Models: Mistral
+We also perform our experiments on Mistral-7B Jiang et al.
+(2023), a popular LLM known for its performance and efficiency.
+Layerrwise distribution of romanized tokens for initial, final and all token generation steps are presented in Figure 19.
+| All tokens | First token | Last token |
+Translation
+Repetition
+Cloze

### v3

@@
 Abstract
-Large Language Models (LLMs) exhibit remarkable multilingual generalization despite being predominantly trained on English-centric corpora.
-A fundamental question arises: how do LLMs achieve such robust multilingual capabilities?
-We take the case of non-Roman script languages, we investigate the role of Romanization—the representation of non-Roman scripts using Roman characters—as a bridge in multilingual processing.
+Large Language Models (LLMs) exhibit strong multilingual performance despite being predominantly trained on English-centric corpora.
+This raises a fundamental question: How do LLMs achieve such multilingual capabilities?
+Focusing on languages written in non-Roman scripts, we investigate the role of Romanization—the representation of non-Roman scripts using Roman characters—as a potential bridge in multilingual processing.
 Using mechanistic interpretability techniques, we analyze next-token generation and find that intermediate layers frequently represent target words in Romanized form before transitioning to native script, a phenomenon we term Latent Romanization.
@@
 These insights contribute to a deeper understanding of multilingual representation in LLMs and highlight the implicit role of Romanization in facilitating language transfer.
+Code and data are available at https://github.com/AI4Bharat/Romanlens.
 RomanLens: The Role of Latent Romanization in Multilinguality in LLMs
@@
 Thanmay Jayakumar1,3,
-Raj Dabre1,3,4,5
+Raj Dabre1,3,4,5444Work done during employment at NICT, Japan.
 Anoop Kunchukuttan1,6,
@@
 (2024) are trained predominantly on English-dominated corpora.
-Yet, they exhibit strong multilingual generalization across diverse languages Huang et al.
+Nonetheless, they exhibit strong multilingual performance across diverse languages Shi et al.
+(2023); Huang et al.
 (2023); Zhao et al.
@@
 This raises a fundamental question: How do LLMs develop such robust multilingual capabilities despite their English-centric training?
-To address this, we adopt a mechanistic interpretability perspective.
-Wendler et al.
-(2024) suggest that LLMs encode multilingual information within a shared, language-agnostic latent space, albeit with an inherent bias toward English due to training data composition and architectural choices.
-Further, Kojima et al.
-(2024) describe distinct phases in multilingual information processing: initial layers map language-specific lexical and syntactic representations to a language-independent semantic space, middle layers maintain this semantic abstraction, and final layers transform these representations into language-specific lexical and syntactic forms.
-In this study, we focus on languages written in non-Roman scripts.
+To address this, prior work by Wendler et al.
+(2024) suggests that LLMs encode multilingual information within a shared, language-agnostic latent space, albeit with an inherent bias toward English due to training data composition and architectural choices.
+Building on this perspective, we investigate a complementary mechanism that may underlie multilingual processing, particularly for languages written in non-Roman scripts.
 We hypothesize that LLMs leverage romanized forms of non-Roman script languages as an intermediate bridge between their language-agnostic concept space and language-specific output representations.
@@
 (2024b).
+Kojima et al.
+(2024) describe distinct phases in multilingual information processing: initial layers map language-specific lexical and syntactic representations to a language-independent semantic space, middle layers maintain this semantic abstraction, and final layers transform these representations into language-specific lexical and syntactic forms.
 Interpretability tools relevant to this work include logit lens Nostalgebraist (2020), tuned lens Belrose et al.
@@
 Direct logit attribution attributes logits to individual attention heads.
-This work focuses on the logit lens (Section 4.1) to investigate whether English-centric decoder only LLMs when prompted in a non-Roman language, processes via romanized latent states before producing native language text.
+This work focuses on the logit lens (Section 4.1) to investigate whether English-centric decoder only LLMs when prompted in a non-Roman script language, processes via romanized latent states before producing native language text.
 Tuned lens is avoided as its training process might obscure the intermediate romanized states by aligning them to final native script outputs, potentially masking the phenomenon under investigation.
@@
 3 Background
-We give a quick background of the transformer’s forward pass as well as the basics of mechanistic interpretability approaches such as logit lens and activation patching which we leverage in this paper.
+We give a quick background of the transformer’s forward pass, romanization and the basics of mechanistic interpretability approaches such as logit lens and activation patching which we leverage in this paper.
 Transformer’s Forward Pass
 Decoder-only transformer models (Vaswani, 2017) employ a residual architecture to process input sequences through multiple layers, producing a sequence of hidden states (latents).
-These latents, whose dimensionality remains the same are updated iteratively across layers through transformer blocks , where indicates the layer index and is the final layer index.
+These latents, whose dimensionality remains the same, are updated iteratively across layers through transformer blocks , where indicates the layer index and is the final layer index.
 For next-token prediction, the final latent is transformed by an unembedding matrix to produce logit scores for vocabulary tokens which are then converted to probabilities via the softmax function (c.f.
 Appendix A).
+3.1 Romanization
+Transliteration is the conversion of text written in one script to another.
+Romanization is a subcategory of transliteration where the target script is English/Latin.
+Within romanization there are multiple romanization schemes available, each based on different considerations.
+One key aspect of romanization schema is if it is lossy or lossless.
+A lossless scheme is required in cases where we have to convert the output back to native script.
+Typically, deterministic transliterations are lossless, whereas natural transliterations are lossy.
+Example:
+The Hindi word for “flower” in Devanagari and its romanization:
+[…]
+Romanization: phool
 3.2 Interpretability Tool: Logit lens
@@
 In the context of activation patching, let denote the language of a word, denote the concept of a word and denote that word.
-Eg: If = cow and = ‘en’ then = ‘cow’.
+For example, if = cow and = ‘en’, then = ‘cow’.
 Similarly = ‘vache’.
-We use 5-shots translation prompts to create paired source and target prompt , with different concept, input language, and output language.
+We use 5-shot translation prompts to create paired source and target prompt , with different concept, input language, and output language.
 Unless otherwise specified, and refer to the output languages of and , respectively.
@@
 RQ1: Do LLMs exhibit latent romanization during multilingual text completion tasks? (Section 4.1)
-RQ2: How does the representation of semantic concepts in LLMs compare between native and romanized scripts of non-Roman languages? (Section 4.2)
+RQ2: How does the representation of semantic concepts in LLMs compare between native and romanized scripts of non-Roman script languages? (Section 4.2)
 RQ3: What are the differences in hidden layer representations when processing the same language in romanized and native scripts? (Section 4.3)
@@
 This is adopted across all analysis setups.
-Appendix B provides a Hindi example of this prompt design, including its English translation and romanized form.
+The prompts are designed around translation, repetition, and cloze tasks, as described below.
 Translation task. We prompt the model to translate a word given five in-context examples.
-Repetition task. We prompt the model to repeat a word in the same language given 5 in-context examples.
+Repetition task. We prompt the model to repeat a word in the same language given five in-context examples.
 Cloze task. We prompt the model to predict the masked word in a sentence given two in-context examples.
+These tasks cover a range of multilingual text completion setups.
+Among these, the repetition task is more syntactic in nature compared to translation and cloze tasks.
+Appendix B provides a Hindi example of prompts across each of the three tasks, along with their English translations and romanized forms.
 4.1 Latent Romanization Analysis
@@
 In contrast, multi-source patching reduces such biases, leading to more robust and generalizable findings.
-In all scenarios from the resulting next token distribution, we compute the probabilities i.e probability of source concept in target language, and i.e probability of target concept in target language.
+In all scenarios from the resulting next token distribution, we compute the probabilities i.e.
+probability of source concept in target language, and i.e.
+probability of target concept in target language.
 We track , i.e., the probability of the concept C occurring in language , by simply summing up the probabilities of all prefixes of and its synonyms in the next-token distribution (c.f.
@@
 To examine the generality of latent romanization, we perform qualitative analyses on five additional languages that use different writing systems: Greek, Ukrainian, Amharic, Hebrew, and Arabic (c.f Appendix H).
+Native | Romanized | |
+Single Source | ||
+Multi Source |
 Language Models:
@@
 (2024) and the translations are romanized using IndicXlit Madhani et al.
-(2023).
+(2023), pypinyin Mozillazg (2024) and Unidecode Šolc (2025).
 All translations were manually validated to ensure data quality.
 6 Results
-We now share our observations about the role of Romanization in LLMs.
-Native | Romanized | |
-Single Source | ||
-Multi Source |
 6.1 Latent romanization
@@
 The translation and cloze tasks function at a semantic level, whereas the repetition task is purely syntactic.
-This means that the repetition task being less complex, the model may reach a decision sooner, potentially in earlier layers and might express its prediction in romanized form in intermediate layers.
-This behavior could be attributed to language-specific neurons, responsible for native script processing, being predominantly concentrated in the initial and final few layers of LLMs Tang et al.
-(2024).
+This means that the repetition task being less complex, the model may reach a decision of what to predict sooner, potentially in earlier layers and might express its prediction in romanized form in intermediate layers.
+This behavior could be attributed to language-specific neurons, responsible for native script processing, being predominantly concentrated in the initial and final few layers of LLMs, Tang et al.
+(2024) leaving the intermediate layers without them.
 6.2 Patching With Romanized Representation vs. Native Representation
@@
 This analysis reveals that LLMs encode semantic concepts similarly regardless of whether the input is in native or romanized script.
-Furthermore, this finding demonstrates that the model achieves comparable levels of language understanding when processing non-Roman languages in their romanized form as in their native script.
+Furthermore, this finding demonstrates that the model achieves comparable levels of language understanding when processing non-Roman script languages in their romanized form as in their native script.
 6.3 Comparing Translations Into Romanized vs. Native Script
@@
 This pattern indicates that when processing non-Roman script languages, the model forms internal representations of target tokens in earlier layers for romanized script compared to native script.
-This trend is consistent across language pairs and models, as shown in Figures 16-18.
+This trend is consistent across language pairs and models (c.f.
+Figures 16-18 in the Appendix).
 This suggests that romanization facilitates faster progression toward language-specific embeddings.
@@
 Romanization acting as a bridge could explain why romanization based script barrier breaking methods like Liu et al.
-(2024); Xhelili et al.
+(2024) and Xhelili et al.
 (2024) work.
@@
 Layerwise analyses reveal that semantic concepts are encoded similarly across native and Romanized inputs, indicating a shared internal representation.
-Moreover, when translating to a Romanized script, target words emerge earlier, highlighting Romanization as a structural link between language-agnostic concepts and language-specific output.
+Moreover, when translating into a Romanized script, target words emerge earlier, highlighting Romanization as a structural link between language-agnostic concepts and language-specific output.
 While our study reveals initial insights into Latent Romanization, future work could focus on applying these findings to develop training strategies that enhance performance across diverse linguistic communities.
@@
 The code and datasets created in this work will be made available under permissible licenses.
-Generative AI systems were only used for assistance purely with the language of the paper, eg: paraphrasing, spell-check, polishing the author’s original content, and for writing boiler-plate code.
+Generative AI systems were only used for assistance purely with the language of the paper, e.g., paraphrasing, spell-check, polishing the author’s original content, and for writing boiler-plate code.
 Acknowledgments
@@
 We keep source concept and target concept distinct to avoid ambiguity when both are expressed in the same target language .
-Cases of token overlap between i.e word representing source concept in target language and i.e word representing target concept in target language and their synonyms are excluded.
+Cases of token overlap between i.e.
+word representing source concept in target language and i.e.
+word representing target concept in target language and their synonyms are excluded.
 Token overlap would cause ambiguity.